OpenStudy (rainbow_rocks03):

PLZ HELP ME I will give you a medal for the BEST response. 70% of a number is 98. What is 90% of that same number? A. 63 B. 68.6 C. 88.2 D. 126

Directrix (directrix):

70% of x = 98 .7*x = 98 x = 140

OpenStudy (amorfide):

if you want 70% of 35 you would do \[\frac{ 70(35) }{ 100 }\] so if you want 70% of an unknown number and you know it is equal to 98 you can write this unknown number using a variable, we will use x so you have \[\frac{ 70x }{100 }\]

OpenStudy (amorfide):

\[\frac{ 70x }{ 100 }=98\]

OpenStudy (amorfide):

work out x, then you will do the same for 90% of the answer you get

Directrix (directrix):

Now, you do this to get the answer: ,9 * 140 = ? Post what you get, okay? @rainbow_rocks03

OpenStudy (amorfide):

so you get 98x100=70x 9800/70=x x=140 90% of 140 is \[\frac{ 90(140) }{ 100 }\]

OpenStudy (rainbow_rocks03):

9*140=1260 @Directrix

OpenStudy (rainbow_rocks03):

never mind.

OpenStudy (amorfide):

yes now divide by 10

Directrix (directrix):

.9*140= 126 @rainbow_rocks03